# Is Productivity in Quantum Programming Equivalent to Expressiveness?

### v1

@@ I Introduction
Programming paradigms describe deliberate choices in code structure and organization, each following specific rules and methods backed by formal rules. Imperative, declarative, functional, and logical paradigms emerged as answers to the question of what form of expressiveness may be more suitable for efficiently solving a problem depending on its nature. Imperative programming focuses on specifying how an algorithm should be implemented. In languages such as FORTRAN, C, C++, Java, or Python, this paradigm enables the construction of algorithms through a sequence of instructions [1]. Object-oriented programming, derived from an imperative base, structures code around objects that encapsulate behaviors and attributes, making it particularly useful for developing complex systems.
@@
Selecting a programming paradigm to solve a problem depends on how well the language form fits the expected functionality within the solution. This choice is not trivial, and it can directly impact code clarity, efficiency, and memory usage. Some problems may be better suited to a specific paradigm, while others may benefit from a combination of multiple paradigms to achieve optimal solutions. With the emergence of quantum computing, the possibility arose that certain problems either difficult or not computable by classical computers could be solved by a quantum computer. This concept was explored by Richard P. Feynman, who proposed that quantum systems could be efficiently simulated only by quantum computers [8]. Despite the evident success of Feynman’s insights, the preliminary formulation of quantum computing centered on computability and tractability rather than programmability, a feature that arises naturally as quantum hardware becomes possible, and later matures.
Although classical programming languages have reached a high level of maturity, they lack the ability to capture the unique properties of quantum computing such as superposition, entanglement, and interference [9]. Instead, most quantum programming languages rely on libraries or domain-specific languages (DSLs) to address these challenges and enable quantum programming. Various software development kits (SDKs), frameworks, and quantum programming libraries are hosted in classical programming languages. Qiskit [10], PennyLane [11], Cirq [12], Qrisp [13], and ProjectQ [14] are examples of tools developed in Python. Quipper, a functional quantum programming language, was impemented in Haskell [15], while quAPL is hosted in APL [16].
@@ II Metrics for Expressiveness Evaluation
Among various dimensions of expressiveness, syntactic expressiveness plays a central role, as it relates to the ability to represent concepts structurally while doing so economically. This has been studied through the analysis of reductions between languages that preserve operational equivalence, providing a means of assessing how abstract and compact a language can be [27, 28]. A key aspect of this analysis is observational equivalence, which holds that two programs or expressions are considered equivalent if, in all possible contexts, they produce the same observable results [29, 28]. Furthermore, syntactic expressiveness entails the fact that transforming constructs from one language to another preserves their observational semantics [28].
@@ II-C Halstead Complexity (HC)
Estimated Time (T)
Approximate development time in seconds:
$T=\frac{E}{18}$ (6)
Halstead suggested that a trained programmer can process approximately 18 mental discriminations per second.
Estimated Number of Errors (B)
Number of potential defects in the program:
$B=\frac{E^{2/3}}{3000}$ (7)
This estimation is useful for evaluating risk and planning testing strategies.
@@ IV-A Selection of Quantum Algorithms
  • Simon’s Algorithm: Implemented using 3 input qubits plus 3 ancillary qubits, with a hidden bitstring ‘”101”‘.
  • Grover’s Search: Evaluated with 3 qubits, performing two iterations, where the marked state was ‘”101”‘.
@@ IV-B Implementation and Code Analysis
We developed a Python script to analyze the source code files and remove non-essential lines, including blank lines and commented lines. To measure Cyclomatic Complexity, we used McCabe’s metric, but applied it at the whole program level rather than to individual functions. Unlike traditional, per-function CC analysis, our implementation evaluates the entire circuit design, which includes the complete quantum algorithm implementation and any oracles used by the algorithm. The CC calculation added one to the total to reflect the number of independent paths within the script. Table I summarizes the constructs considered in each language to account for control flow complexity within CC.
@@
Qiskit, Qrisp, Cirq (Python) if, elif, for, while, except, with, assert, list, set, dict, and, or
@@
Qmod if, else, repeat, within, apply, lambda, and, or
Q# if, elif, for, ApplyToEachA, try, catch, repeat, until
OpenQASM 2.0 if, else
@@ IV-C Scope and Limitations
We chose four fundamental algorithms were used for evaluation in order to consistently build the evaluation methodology with compact yet meaningful examples. The proposed methodology may be similarly expanded to include more algorithms in future exercises. We are aware that our choice of algorithms only consists of the most simple ones, which may bias the results to a subset of possible algorithmic patterns.
@@ V-A Lines of Code Analysis
Language Deutsch-Jozsa Bernstain-Vazirani Simon Grover
@@
Cirq, Qiskit, and Qrisp share similar results. These three languages have closely matching LOC across all implementations. Since they share Python as their host language, they likely follow similar code structures and rely on comparable libraries that optimize algorithm implementations uniformly. Q# consistently requires more lines of code. Every algorithm written in Q# shows relatively high LOC, indicating that its modular design or strong typing may demand more explicit code compared to other languages.
Results differ substantially between standalone quantum programming languages. quAPL displays significant variation in LOC. While it requires a high number of lines for Grover (44) and Deutsch-Jozsa (35), it remains extremely compact for Bernstein-Vazirani (11). This variation suggests that quAPL’s implementation is optimized for some algorithms but less efficient for others. Qmod is the only language where Bernstein-Vazirani surpasses Deutsch-Jozsa in LOC. In every other language, Deutsch-Jozsa requires more lines than Vazirani. However, in Qmod, Bernstain-Vazirani (25) exceeds Deutsch-Jozsa (29), which suggests structural differences in how Qmod implements these algorithms.
Q# and quAPL show the most extreme differences in LOC. Q# consistently requires a high number of lines across all algorithms, while quAPL displays the widest range of LOC. These results indicate that both languages define different routes to code expressiveness and structure. Despite being a crude metric, average LOC per language are suggestive of differences in overall verbosity across languages independent of specific algorithms (Fig. 2). Qmod exhibits the lowest average number of lines of code. This language maintains the most compact implementation of the evaluated algorithms, suggesting the presence of efficient abstractions to reduce code verbosity. By contrast, Cirq, quAPL, and Qiskit have similar average LOC values. These three languages fall into an intermediate position in terms of verbosity, sharing a similar level of expressiveness when implementing quantum algorithms.
@@ V-B Cyclomatic Complexity Analysis
Cirq 21 7 18 36
quAPL 3 10 5 12
Qiskit 16 5 14 39
Qrisp 10 5 10 37
Qmod 9 8 4 4
Q# 6 7 7 15
@@
Grover remains the algorithm with the highest cyclomatic complexity in most of the evaluated languages. Qiskit, with a score of 39, Qrisp with 37, and Cirq with 36, have the most complex implementations in terms of CC. These values suggest that these languages require a significant number of control structures to implement Grover, which aligns with their behavior in terms of lines of code. However, Qmod behaves unusually, with a score of only 4 for Grover, the lowest among all languages and algorithms. This result is particularly notable because its cyclomatic complexity for Deutsch-Jozsa and Vazirani is higher, with values of 9. This suggests Qmod includes abstractions in its library that privilege efficient representation of transformations appearing in the algorithm.
@@
Cirq and Qiskit follow similar trends across all algorithms, with complexity scores that remain closely aligned. This finding reinforces the idea that both languages, by sharing Python as their host language, may adopt similar strategies for structuring code and implementing quantum algorithms. Their complexity scores in Deutsch-Jozsa (21 for Cirq and 16 for Qiskit), Bernstein-Vazirani (7 and 5, respectively), and Simon (18 and 14) reflect this trend, showing only small differences in each case. quAPL has the lowest cyclomatic complexity across all evaluated algorithms. Its scores range from 3 to 12, with particularly low values in Deutsch-Jozsa (3) and Simon (5). This is expected, since its host language (APL) usually results in few control structures for direct statements.
Qrisp and Q# fall into an intermediate range in terms of cyclomatic complexity. Q# maintains relatively low cyclomatic complexity across all algorithms, with scores never exceeding 15, suggesting that its modular design may contribute to a simpler code structure with powerful enough control flow. Qmod behaves particularly atypically in Simon and Grover, showing the lowest values among all languages. In Simon, it scores only 4, while in Grover it also scores 4, contrasting sharply with the significantly higher values found in other languages such as Cirq (18 for Simon and 36 for Grover) or Qiskit (14 and 39, respectively). Once again Qmod appears to optimize for frequently used quantum patterns.
@@
Overall, cyclomatic complexity varies significantly across languages, showing patterns consistent with previous findings on lines of code. Grover remains the most complex algorithm in most cases, with the exception of Qmod, which notably breaks this trend. Cirq and Qiskit maintain similarities in their control structures, while quAPL stands out for its extreme simplicity across all algorithms. Qrisp and Q# remain at an intermediate level, with Q# showing particularly low values. These results reinforce the expectation that the choice of quantum programming language impacts not only the amount of written code but also the structural complexity of the implementation. These trends are further validated by Figure 4, which shows the mean CC for all quantum programming languages. The highest complexity levels correspond to Python hosted languages, particularly Cirq and Qiskit, while standalone languages remain more compact and manageable in terms of control flow complexity.
@@ V-C Complexity Metrics Analysis
Given the richness of the Halstead Complexity metrics, we decided to evaluate the relationship between Lines of Code (LOC), Cyclomatic Complexity (CC), Effort, Volume, Difficulty, and Vocabulary, following the definitions by Halstead and McCabe (Fig.  5). Fig.  6 summarizes the findings reported below.
@@
We found a clear separation between Python-hosted languages and stand-alone alternatives. This points to a programming style in which the cost of classical syntax for detailed control flow is overimposed upon quantum concerns. Q# stands out as the language with the highest LOC but a relatively controlled CC, indicative of lower syntactic cost to achieve similar results compared to Cirq or Qiskit. Qmod once again breaks the trend with the lowest values in both metrics. This language has both the fewest lines of code and the lowest cyclomatic complexity, indicating a highly optimized implementation that minimizes the amount of code and control decisions required to execute the algorithms. Qrisp holds an intermediate position.
@@
While quAPL may be concise, it shares a similar place with Cirq in having the highest effort values. Q# seems to follow a somewhat linear trend between LOC and effort alongise Qiskit, Qmod and Qrisp . Qmod remains at the lower end for both LOC and effort, suggesting minimal cognitive load in proportion to code volume.
@@
Qmod and Cirq sit at opposite extremes of the CC-effort spectrum. Q# and quAPL cluster together, so as Qrisp and Qiskit. We found this surprising due to the differences and similarities shared across languages. quAPL and Q# are concise and intellectually expensive, which is consistent with what we expect from APL as well as the underlying cost of guarantees provided by Q#. Qrist and Qiskit are similar as expected, while design choices in Cirq may suggest a syntax closer to quantum circuit fundamentals.
@@
All evaluated languages fit a linear trend between volume (information content) vs effort. Qmod remains as the most compact and easily interpretable in our evaluation set. Despite being compact quAPL appears to force program writers to unpack large information contents at high mental effort; surprisingly, Cirq does as well despite belonging to a different programming paradigm. The remaining languages appear to show metrics determined by the absence (Qiskit, Qrisp) or presence (Q#) of static type checking.
V-C1 Difficulty vs Effort
@@ V-C2 LOC vs Vocabulary
n1 (Unique Operators) 23.75 28.25 29.75 19.75 13.00 18.75
n2 (Unique Operands) 13.25 13.75 17.75 13.50 8.75 10.50
N1 (Total Operators) 117.00 86.50 115.75 50.25 27.00 45.75
N2 (Total Operands) 61.50 52.75 62.25 60.25 32.75 50.00
Vocabulary 37.00 42.00 47.50 33.25 21.75 29.25
Length 178.50 139.25 178.00 110.50 59.75 95.75
Volume 938.61 758.65 1001.72 569.56 264.66 475.02
Difficulty 55.21 55.43 47.85 44.23 24.76 47.98
Effort 59390.94 47715.10 57809.82 29898.75 6629.53 28814.38
@@ VI Discussion
Overall, the results show a consistent correlation between LOC and CC in Python-hosted languages, while Qmod and quAPL diverge from this pattern. Halstead metrics further differentiate languages by capturing lexical complexity and cognitive effort. The complexity introduced by static typing and a classical-quantum syntax in Q# breaks a few expectations. Q# consistently shows the highest average number of lines of code, confirming its verbose nature across all evaluated algorithms. Nevertheless, its cyclomatic complexity remains moderate, indicating that the language promotes explicit and detailed syntax while limiting the use of branching control structures.
Cirq and Qiskit show strong alignment in both LOC and CC metrics. These languages share a common host environment (Python), which may explain their structural similarity and predictable behavior. Their results display similar patterns across the board, reinforcing the idea that they follow a shared modular and syntactic design. Qrisp consistently occupies a middle ground across our experiments. Its LOC and CC values are lower than those of Cirq or Qiskit but higher than Qmod. This balance between conciseness and structural control suggests a design philosophy that avoids extremes in verbosity or abstraction.
Qmod deviates from these patterns. It records the lowest LOC and CC values, along with minimal effort and difficulty scores based on Halstead’s definitions. These results indicate that Qmod supports minimalistic representations of quantum algorithms, optimizing both code size and structural simplicity. In contrast to other languages, Qmod maintains uniform complexity levels across all evaluated algorithms, including Grover. quAPL also exhibits an atypical profile. Although it maintains low cyclomatic complexity, its effort and difficulty metrics remain high. This combination suggests that, despite its structural simplicity, quAPL implementations require more cognitive processing due to symbolic density and lexical variety. Notably, both Qmod and quAPL allow for functional programming constructs, which may contribute to improved results in certain metrics by reducing control flow branching without sacrificing expressiveness.

### v2

@@ I Introduction
Programming paradigms describe deliberate choices in code structure and organization, each following specific rules and methods backed by formal rules. Imperative, declarative, functional, and logical paradigms emerged as answers to the question of what form of expressiveness may be more suitable for efficiently solving a problem depending on its nature. Imperative programming focuses on specifying how an algorithm should be implemented. In languages such as FORTRAN, C, C++, Java, or Python, this paradigm enables the construction of algorithms through a sequence of instructions [1]. Object-oriented programming, with an imperative origin, structures code around objects that encapsulate behaviors and attributes, making it particularly useful for developing complex systems.
@@
Selecting a programming paradigm to solve a problem depends on how well the language form fits the expected functionality within the solution. This choice is not trivial, and it can directly impact code clarity, efficiency, and memory usage. Some problems may be better suited to a specific paradigm, while others may benefit from a combination of multiple paradigms to achieve optimal solutions. With the emergence of quantum computing, the possibility arose that certain problems either difficult or not computable by classical computers could be solved by a quantum computer. This concept was explored by Richard P. Feynman, who proposed that quantum systems could be efficiently simulated only by quantum computers [8]. Despite the evident success of these insights, the preliminary formulation of quantum computing centered on computability and tractability rather than programmability, a feature that arises naturally as quantum hardware becomes possible, and later matures.
Although classical programming languages have reached a high level of maturity, they lack the ability to capture the unique properties of quantum computing such as superposition, entanglement, and interference [9]. Instead, most quantum programming languages rely on libraries or domain-specific languages (DSLs) to address these challenges and enable quantum programming. Various software development kits (SDKs), frameworks, and quantum programming libraries are hosted in classical programming languages. Qiskit [10], PennyLane [11], Cirq [12], Qrisp [13], and ProjectQ [14] are examples of tools developed in Python. Quipper, a functional quantum programming language, was implemented in Haskell [15], while quAPL is hosted in APL [16].
@@ II Metrics for Expressiveness Evaluation
Among various dimensions of expressiveness, syntactic expressiveness plays a central role, as it relates to the ability to represent concepts structurally while doing so economically. This has been studied through the analysis of reductions between languages that preserve operational equivalence, providing a means of assessing how abstract and compact a language can be [27, 28]. A key aspect of this analysis is observational equivalence, which holds that two programs or expressions are considered equivalent if, in all possible contexts, they produce the same observable results [29, 28]. Furthermore, syntactic expressiveness entails transforming constructs from one language to another while preserving their observational semantics [28].
@@ IV-A Selection of Quantum Algorithms
  • Simon’s Algorithm: Implemented using 3 input qubits plus 3 ancillary qubits, with a hidden bitstring ‘101‘.
  • Grover’s Search: Evaluated with 3 qubits, performing two iterations, where the marked state was ‘101‘.
@@ IV-B Implementation and Code Analysis
We developed a Python project to analyze the source code files and remove non-essential lines, including blank lines and commented lines. To measure Cyclomatic Complexity, we used McCabe’s metric, but applied it at the whole program level rather than to individual functions. Unlike traditional, per-function CC analysis, our implementation evaluates the entire circuit design, which includes the complete quantum algorithm implementation and any oracles used by the algorithm. The CC calculation added one to the total to reflect the number of independent paths within the script. Table I summarizes the constructs considered in each language to account for control flow complexity within CC.
@@
Qiskit, Qrisp, Cirq (Python) if, elif, for, while, except, with, assert, list, set, dict, and, or, on_each
@@
Qmod if, else, repeat, within, apply, lambda, and, or, on_each
Q# if, elif, for, ApplyToEachA,MeasureEachZ, try, catch, repeat, until
@@ IV-C Scope and Limitations
We chose four fundamental algorithms in order to consistently build the evaluation methodology with compact yet meaningful examples. The proposed methodology may be similarly expanded to include more algorithms in future exercises. We are aware that our choice of algorithms only consists of the most simple ones, which may bias the results to a subset of possible algorithmic patterns.
@@ V-A Lines of Code Analysis
Language Deutsch-Jozsa Bernstein-Vazirani Simon Grover
@@
Cirq, Qiskit, and Qrisp share similar results. These three languages have closely matching LOC across all implementations. Since they share Python as their host language, they likely follow similar code structures and rely on comparable libraries that optimize algorithm implementations uniformly.
Q# consistently requires more lines of code than any other language. For example, it reaches 33 LOC for Deutsch-Jozsa and 56 for Grover. This suggests that its modular architecture and strong static typing introduce additional verbosity in algorithm construction, especially in oracle definitions and measurement logic.
quAPL, also demonstrates consistently high LOC across algorithms. Despite the syntactic brevity of APL, quAPL records 35 LOC for Deutsch-Jozsa, 20 for Bernstein-Vazirani, and 44 for Grover. This indicates that while quAPL may use dense expressions, its functional style still demands structural clarity that translates into longer code.
Qmod stands out for its highly compact implementation of Grover’s algorithm (9 LOC), which is noticeably lower than its implementations of Bernstein-Vazirani (25 LOC) and Deutsch-Jozsa (29 LOC). This difference suggests that Qmod includes specific abstractions optimized for Grover, whereas other algorithms may still require more explicit or verbose logic.
These results indicate that both languages define different routes to code expressiveness and structure. Despite being a crude metric, average LOC per language are suggestive of differences in overall verbosity across languages independent of specific algorithms (Fig. 2). Qmod exhibits the lowest average number of lines of code. This language maintains the most compact implementation of the evaluated algorithms, suggesting the presence of efficient abstractions to reduce code verbosity. By contrast, Qrisp, Qiskit, Cirq, and quAPL have similar average LOC values. These three languages fall into an intermediate position in terms of verbosity, sharing a similar level of expressiveness when implementing quantum algorithms.
@@ V-B Cyclomatic Complexity Analysis
Cirq 6 6 6 11
quAPL 3 4 5 12
Qiskit 2 3 3 8
Qrisp 2 3 3 8
Qmod 5 5 2 2
Q# 5 6 6 12
@@
Grover remains the algorithm with the highest cyclomatic complexity in most of the evaluated languages. Q# and quAPL with a score of 12, and Cirq with 11, have the most complex implementations in terms of CC. These values suggest that these languages require a significant number of control structures to implement Grover, which aligns with their behavior in terms of lines of code. However, Qmod behaves unusually, with a score of only 2 for Grover and Simon, the lowest among all languages. This result is particularly notable because its cyclomatic complexity for Deutsch-Jozsa and Bernstein-Vazirani is higher, with values of 5. This suggests Qmod includes abstractions in its library that privilege efficient representation of transformations appearing in the algorithm.
@@
Qrisp and Qiskit follow similar trends across all algorithms, with complexity scores that remain closely aligned. This finding reinforces the idea that both languages, by sharing Python as their host language, may adopt similar strategies for structuring code and implementing quantum algorithms. The complexity in their scores for Deutsch-Jozsa (2 for both), Bernstein-Vazirani (3 each), Simon (3 each), and Grover (8 for Qiskit, 8 for Qrisp) reflect this consistency. quAPL scores range from 3 to 12, with particularly intermediate values in Bernstein-Vazirani and Simon (5). Notably, quAPL presents some of the highest individual cyclomatic complexity scores, particularly in Grover (12). While its overall complexity is not consistently the highest across all algorithms, these peaks suggest that APL’s terse and expressive style may lead to dense implementations that involve non-trivial control flows in certain cases. Cirq and Q# exhibit the highest average cyclomatic complexity among all evaluated languages, both with mean scores of 7.25 across the four algorithms. This suggests that, while both languages offer rich control-flow mechanisms and abstraction features, their implementations may involve more branching and conditional logic.
@@
Overall, cyclomatic complexity varies significantly across languages, showing patterns consistent with previous findings on lines of code. Grover remains the most complex algorithm in most cases, with the exception of Qmod, which notably breaks this trend. Cirq and Q# exhibit the highest average complexity, reflecting more extensive use of control structures and branching. In contrast, Qiskit and Qrisp maintain low and nearly identical complexity scores, suggesting streamlined implementations possibly influenced by Python’s expressive syntax. quAPL shows moderate to high complexity in some algorithms, particularly Grover. These results reinforce the idea that the choice of quantum programming language affects not only the size of the codebase but also the structural intricacy of the implementation. These trends are further supported by Figure 4, which illustrates the average cyclomatic complexity per language. The highest complexity levels are observed in Cirq and Q#, while Qmod stands out for offering the most compact control flow across all evaluated cases.
@@ V-C Complexity Metrics Analysis
Given the richness of the Halstead Complexity metrics, we decided to evaluate the relationship between Lines of Code (LOC), Cyclomatic Complexity (CC), Effort, Volume, Difficulty, and Vocabulary, following the definitions by Halstead and McCabe (Fig.  5). Additionally, Fig.  6 summarizes the findings reported below.
@@
We found some degree of separation between languages that favor high-level abstractions and those that rely on more explicit control structures. Q# and Circ stands out with the highest lines of code and cyclomatic complexity, reflecting a programming model that emphasizes explicit quantum operation declarations and strict type enforcement. quAPL also exhibits high LOC and CC, nearly on par with Cirq, despite being based on APL, a language known for terse syntax.
In contrast, Qiskit and Qrisp cluster together in the lower-left region of the plot, displaying both low LOC and low CC. This pattern can be attributed to the use of array-based abstractions and internal iteration mechanisms. For instance, in Qrisp, operations like qrisp.h(input) or qrisp.cx(a, b) are implicitly broadcasted over quantum registers, eliminating the need for explicit for loops in most quantum algorithms. Similarly, Qiskit often utilizes Pythonic range-based constructs and functional programming patterns that reduce syntactic overhead. These abstractions help streamline circuit construction and reduce branching, which directly impacts cyclomatic complexity. Finally, Qmod consistently appears as the most compact language in both metrics, indicating a declarative and minimalistic approach that simplifies both the structural and syntactic layers of quantum program design.
@@
quAPL shares a similar place with Cirq in having the highest effort values. Q# seems to follow a somewhat linear trend between LOC and effort alongise Qiskit, Qmod and Qrisp . Qmod remains at the lower end for both LOC and effort, suggesting minimal cognitive load in proportion to code volume.
@@
quAPL, Cirq, and Q# occupy the upper end of both cyclomatic complexity and Halstead effort, suggesting high structural and cognitive demands. quAPL reaches the highest effort overall, likely due to the dense lexical nature of APL, which, despite its brevity, encodes significant semantic complexity. Cirq follows with similarly high scores, reflecting its low-level, imperative design. At the opposite end, Qmod shows the lowest complexity and effort, reinforcing its compact style. Qiskit and Qrisp form a middle ground: both maintain low CC, but moderate effort, likely due to their use of expressive abstractions over Python’s syntax, which reduce control structures but still entail cognitive overhead.
@@
All evaluated languages fit a linear trend between volume (information content) vs effort. Qmod remains as the most compact and easily interpretable in our evaluation set. Despite being symbolic compact quAPL appears to force program writers to unpack large information contents at high mental effort; surprisingly, Cirq does as well despite belonging to a different programming paradigm. The remaining languages appear to show metrics determined by the absence (Qiskit, Qrisp) or presence (Q#) of static type checking.
Difficulty vs Effort
@@
n1 (Unique Operators) 24.00 29.25 28.50 20.75 15.50 20.00
n2 (Unique Operands) 13.25 13.75 19.00 13.50 8.75 10.50
N1 (Total Operators) 127.50 105.75 152.50 79.75 59.25 71.25
N2 (Total Operands) 61.50 52.75 65.75 60.25 32.75 50.00
Vocabulary 37.25 43.00 47.50 34.25 24.25 30.50
Length 189.00 158.50 218.55 140.00 92.00 121.25
Volume 995.22 868.27 1234.53 727.26 420.47 608.55
Difficulty 55.71 57.34 48.40 46.41 29.76 50.76
Effort 63116.99 56107.02 75780.32 39869.05 12781.16 38853.50
@@ VI Discussion
Overall, the results show varying relationships between LOC and CC across languages. While Qiskit and Qrisp maintain both metrics at low levels, likely due to their use of high-level abstractions, Cirq exhibits high cyclomatic complexity despite moderate LOC, indicating a more granular and imperative coding style, even though it is also based on Python. Q# sits at the upper end of both dimensions, reflecting verbose and structurally complex implementations. Halstead metrics further distinguish languages by capturing lexical density and cognitive effort. Q#
Qrisp and Qiskit show strong alignment in both LOC and CC metrics. These languages share a common host environment (Python), which may explain their structural similarity and predictable behavior. Their results display similar patterns across the board, reinforcing the idea that they follow a shared modular and syntactic design. Qrisp consistently occupies a middle ground across our experiments. Its LOC and CC values are lower than those of Cirq or Qiskit but higher than Qmod. This balance between conciseness and structural control suggests a design philosophy that avoids extremes in verbosity or abstraction. Qmod records the lowest LOC and CC values, along with minimal effort and difficulty scores based on Halstead’s definitions.
Qmod stands out for its concise and declarative syntax, which enables the expression of complex logic using compact functional constructs. As illustrated by its use of inline lambda functions and nested logical operators, for example, in Grover’s algorithm, Qmod avoids explicit control structures in favor of composable quantum primitives. This design contributes directly to its minimal cyclomatic complexity and low Halstead effort across all evaluated algorithms. Its support for functional programming appears to streamline circuit construction while reducing structural and syntactic overhead, making it a compelling language for efficient quantum programming.
On the other hand, quAPL exhibits an atypical profile. Despite its syntactically concise nature, it shows high cyclomatic complexity, effort, and difficulty scores. This suggests that quAPL implementations, while brief in appearance, demand substantial cognitive processing due to symbolic density and lexical variety. Notably, both Qmod and quAPL support functional programming constructs, which may contribute to reduced reliance on explicit branching. However, only Qmod consistently translates these constructs into lower structural and lexical complexity, while quAPL retains a high semantic load.